15(N+15)=0
SOLVING EQUATIONS BY FACTORING

Answer by bacos(9)

Simplifying
15(N + 15) = 0
Reorder the terms:
15(15 + N) = 0
(15 * 15 + N * 15) = 0
(225 + 15N) = 0
Solving
225 + 15N = 0
Solving for variable 'N'.
Move all terms containing N to the left, all other terms to the right.
Add '-225' to each side of the equation.
225 + -225 + 15N = 0 + -225
Combine like terms: 225 + -225 = 0
0 + 15N = 0 + -225
15N = 0 + -225
Combine like terms: 0 + -225 = -225
15N = -225
Divide each side by '15'.
N = -15
Simplifying
N = -15